troy2225:

In the equation 6x – 2 = –4x + 2, Spencer claims that the first step is to add 4x to both sides. Jeremiah claims that the first step is to subtract 6x from both sides. Who is correct? Explain.

Mercury:

Please do not copy-paste my answer. Since we want all the x-terms on one side, both of their approaches work Using Jeremiah's logic: 6x – 2 = –4x + 2 subtracting 6x from both sides -2 = -10x + 2 Using Spencer logic: 6x – 2 = –4x + 2 subtracting 4x to both sides 10x - 2 = 2